There is 5 grams of C^2H^6 in a combustion reaction , how many grams of each product is produced


1
Expert's answer
2021-04-29T07:18:07-0400

2C2H6 + 7O2 → 4(CO2) + 6H2O


Moles of C2H6"= \frac{5}{30.07}=0.166mol"




Mass of CO2= 2 × 0.166× 44.01= 14.61grams


Mass of H2O = 3 × 0.166 × 18.02 = 8.97 grams